OpenStudy (anonymous):

Equation... Thought it was no problem as I was taught by heather ... This is hard for me .... X(4x-3)=(2x-1)^2 ? Explain :(

OpenStudy (anonymous):

exand, try to solve for x by getting a quadratic equation

OpenStudy (anonymous):

X(4x-3)=(2x-1)^2 (-12x)=(-2x)^2 Add 12 both sides X=14x^2. ? And ?

OpenStudy (anonymous):

No, okay lets look at the left side only you have x(4x-3) okay? Now your going to FOIL. so you will have 4x^2 and -3x do you see how i get that?

OpenStudy (anonymous):

Yuh ok...

OpenStudy (anonymous):

Do you know how to FOIL? or expand?

OpenStudy (anonymous):

anyway your going to do the same thing to the right side so what do you get?

OpenStudy (anonymous):

Yuh

OpenStudy (anonymous):

Sorry for the late reply . Wifi on my iPad f'd up